March 12 in GB is the plan. 4 rooms Jianth UC-b will do res
March Testing Status
Mode 1 working and tested at high clock rate more testing ok for march
filters will be ~1.25 GHz bandwidth (1250 MHz per roach)
Roach boards from UC-B for March test can be shipped this week.
Use one Roach for RFI sweeper and 6 others for observations.
Computers for test
We'll have one computer for each ROACH (7) but they do not have enough to store data locally (we will add storage (two disks: 2 or 3TB?). UC-B will bring a filer computer if needed if NRAO 'pork' is unavailable. All roaches and computers will attach to a 10 gbe switch. NRAO has the machines and the 10 gbe cards for them. Disk drives will be shipped directly to NRAO.
We could connect the roaches directly to the CPU's, and then connect all the CPU's to the filer via 1Gbit. UC-B will check that filer can handle aggregate data rate (2Mbit/sec to disk). Data will be stored in SDFITS and converted to PSRFITS offline.
waiting for IF PCBs - arrived, will build two for RFI checkout then build balance.
Delivery of the borrowed ADCs in Green Bank by the 26th preferred.
NRAO Modification Request (MR) status: Noted as'VEGAS'in NRAO Plan of Record
Software integration tasks
ADC sampling (I think Patrick was waiting on a Berkley dependency for this)
Mark W will add them and update BOF file this week
Whether Ron had supplied info to Patrick about the DDL for the Coordinator
No, Patrick will check
Roach switching signal tables
Code for setup in place, needs verifying
FW update for Roaches (again, Berkley dependency)
IF/LO setup software
UC-B can send some snap blocks for testing, Melinda working on Config.
Needs most of the system in place to test
Integrated testing of VEGAS manager with other GBT software and checking the data with GBTIDL.